OpenStudy (anonymous):

Find the slope, if it exists, of the following line: 3/4y=-7/9x

OpenStudy (anonymous):

-28/27

OpenStudy (anonymous):

multiply both sides by 4/3 (4/3)3/4y=(4/3)(-7/9)x or y = -28/27x The slope is -28/27

OpenStudy (anonymous):

multiply both sides by 4/3. you get -7/9 times 4/3 which is the answer above.

OpenStudy (anonymous):

Thank you very much!